Solve by completing the square.
x^2 – 10x + 24 = 0

Answer by Nate(3500)

x^2 – 10x + 24 = 0
x^2 – 10x = -24
(x - 5)^2 = -24 + 25 = 1
x - 5 = +- 1
x = 5 +- 1
x = 6 and x = 4